Consider the equation: +5 0 A) First, use the completing the square process to write this equation in the form (z + D-orD)
Answer #1


2 2 2
Answer #2

SOLUTION :


A.


x^2 + 5x = 0


Converting to completed square form :


=> (x + 5/2)^ - (5/2)^2 =  0

=> (x + 5/2)^2 = 25/4  


So , x^2 + 5x is equivalent to :


( x + 5/2)^2 = (5/2)^2  (ANSWER).


B.


(x + 5/2)^2 = (5/2)^2 

Taking square root on both sides :

=> x + 5/2 = +/- 5/2

=> x = 0 , - 5 (ANSWER) 

answered by: Tulsiram Garg